Find the present value of $700 due in the future under each of the following conditions. Do not round intermediate calculations. Round your answers to the nearest cent.

  1. 9% nominal rate, semiannual compounding, discounted back 5 years.

    $  

  2. 9% nominal rate, quarterly compounding, discounted back 5 years.

    $  

  3. 9% nominal rate, monthly compounding, discounted back 1 year.

    $  
Answer #1

FV =700

a. Semi annual Rate =9%/2 =4.5%
Number of Periods =5*2 =10
PV =FV/(1+r)^n =700/(1+4.5%)^10 =450.75

b. Quarterly Rate =9%/4 =2.25%
Number of Periods =5*4 =20
PV =FV/(1+r)^n =700/(1+2.25%)^20 =448.57

c. Monthy Rate =9%/12 =0.75%
Number of Periods =1*12 =12
PV =FV/(1+r)^n =700/(1+0.75%)^12 =639.97